# Purification Before Fusion: Toward Mask-Free Speech Enhancement for Robust Audio-Visual Speech Recognition

### v1

@@ Abstract
Audio-visual speech recognition (AVSR) typically improves recognition accuracy in noisy environments by integrating noise-immune visual cues with audio signals. Nevertheless, high-noise audio inputs are prone to introducing adverse interference into the feature fusion process. To mitigate this, recent AVSR methods often adopt mask-based strategies to filter audio noise during feature interaction and fusion, yet such methods risk discarding semantically relevant information alongside noise. In this work, we propose an end-to-end noise-robust AVSR framework coupled with speech enhancement, eliminating the need for explicit noise mask generation. This framework leverages a Conformer-based bottleneck fusion module to implicitly refine noisy audio features with video assistance. By reducing modality redundancy and enhancing inter-modal interactions, our method preserves speech semantic integrity to achieve robust recognition performance. Experimental evaluations on the public LRS3 benchmark suggest that our method outperforms prior advanced mask-based baselines under noisy conditions.
@@ 1 Introduction
Most contemporary AVSR models employ cross-modal attention mechanisms to facilitate effective audio-visual interaction [6, 7, 8, 9, 10]. Compared to early strategies such as direct feature concatenation, cross-attention achieves dynamic alignment and richer multimodal integration. But, when audio inputs are severely corrupted by noise, the interaction process is prone to disruption. Noisy audio features can introduce irrelevant or misleading information, forcing the model to undertake the entangled tasks of both implicitly “denoising” the input and “extracting” critical speech-related information simultaneously. This dual burden might overwhelm the cross-modal interaction module, leading to suboptimal feature fusion. To alleviate the negative effects of noise, recent studies have proposed custom-designed masking networks to generate noise-reduction masks, which explicitly suppress irrelevant noise components in audio features [4, 11, 12]. These methods aim to suppress noise by selectively focusing on relevant audio information during fusion. Nevertheless, they are typically driven solely by the final AVSR objective, which cannot guarantee the preservation of semantic integrity during this lossy noise-suppression process.
@@ 2 Proposed Method
Fig. 1 illustrates the proposed noise-robust AVSR framework, which processes video and noisy audio inputs to achieve speech recognition without relying on explicit masks. The framework integrates a primary AVSR model with a dedicated speech enhancement (denoising) module that reconstructs cleaner audio representations. The enhancement module is placed after audio feature extraction and before cross-modal fusion, and it is jointly trained end-to-end with the AVSR model. The refined audio features are then passed through a cross-modal encoder and fused with the visual features, enabling deeper and more reliable speech understanding.
@@ 2.2 Audio-Visual Bottleneck Conformer
where $\theta_{v}$ and $\theta_{a}$ are the parameters of the visual Conformer and the audio Conformer, respectively. $\|$” denotes the concatenation operation of the latent tokens. Note that the convolution module within the Conformer block is only used for modality tokens excluding the bottleneck tokens. The iterative bottleneck attention computation across $\mathbf{h}_{a}$ and $\mathbf{h}_{v}$ ultimately yields compressed representations $\mathbf{z}_{v}=\mathbf{h}_{v}^{L}$ and $\mathbf{z}_{a}=\mathbf{h}_{a}^{L}$. Since all cross-modal attention flow must pass through these bottleneck tokens, such tight “fusion” bottlenecks force the model to condense modality-specific information and share only essential content. This design allows the visual modality to guide the purification of noisy audio features in a computationally efficient manner, which reduces cross-modal attention computations from $\mathcal{O}((N_{a}+N_{v})^{2})$ to $\mathcal{O}((K+N_{a})^{2})+\mathcal{O}((K+N_{v})^{2})$.
@@ 3.2 Experimental Settings
Testing Setup. We utilize the model averaged over the last 10 checkpoints for evaluation following [23]. During inference, decoding is performed with a left-to-right beam search of width 40 on the Transformer decoder, and following [27], we exploit an off-the-shelf pre-trained GPT-2 language model [28] for beam rescoring. Following prior literature [1, 2], we report word error rate (WER) on LRS3 under both clean and noisy conditions with varying SNR levels. A lower WER indicates better recognition performance, while a lower SNR corresponds to a higher noise level.
@@
Table 3: WER (%) comparisons of different methods under clean audio and overlapped speech (SNR=-5dB) inputs, with or without video assistance.
@@ 3.3.1 Influence of the number of bottleneck tokens
The bottleneck tokens serve a key role in our AVBC fusion structure. We thus conduct experiments to evaluate how their quantity influences the model’s final prediction performance under -5dB babble noise. We keep the number of bottleneck tokens to be much smaller than the total latent units per modality, explicitly ensuring only necessary information resides in these bottleneck tokens. Specifically, we set the number of tokens to 0, 1, 2, 4, 8, and 16 respectively for comparison, where “0” means that the two modalities skip the bottleneck and directly engage in cross-attention. As depicted in Fig. 2, we can observe that the model achieves relatively optimal performance when the number of tokens is set to 4. The results imply that too few tokens may hinder sufficient exchange of essential inter-modal information, while too many could compromise the model’s ability to prioritize only essential content transmission.
@@ 3.3.3 Noise robustness evaluation
To investigate the noise robustness of our model against prior competitive baselines, we perform comparative experiments across different acoustic conditions: clean audio and noisy settings with an SNR range of -5 to 15 dB. As shown in Table 2, our proposed method outperforms these advanced noise-robust methods that require explicit noise-reduction masks (e.g., AV-RelScore [12], which incorporates a reliability scoring module to suppress corrupted modal representations) by achieving a lower average WER. Particularly, as SNR level decreases, the performance gap between our method and these baselines widens. Moreover, our full model reduces WER by 1.7% compared to the ablation variant without speech feature enhancement objectives. The results suggest that our method can effectively suppress noisy audio representations without relying on explicit noise mask generation.
@@ 3.3.4 Experiments on various input conditions
Since cross-modal attention operations are decoupled from the modal encoders, our AVSR framework generalizes to single-modality input sequences, aligning with [9]. To examine the model’s robustness under diverse input conditions, we further conduct evaluations covering clean audio and overlapped speech, each assessed with and without video input. Overlapped speech is generated by randomly selecting a sample from the test set that is distinct from the input audio and overlapping it with the input audio at an SNR of -5dB. During testing, the two distinct speech signals are played simultaneously, while the input video only provides visual cues for the target speech. Notably, the bottleneck tokens within the AVBC still interact with the audio sequence even when video input is disabled. As shown in Table 3, the presence or absence of video input exerts minimal influence on the final performance under clean audio conditions. However, when overlapped speech is used as the audio input, the visual modality becomes crucial for effectively “selecting” the target speech from the mixed speech. Furthermore, our model (without speech enhancement) performs better than the baseline in [9] under overlapped speech scenarios. This advantage likely stems from the bottleneck tokens preserving modality-shared features, which serve as complementary cues for the corrupted audio.
@@ 4 Conclusion
In this paper, we propose a noise-robust AVSR architecture that integrates AVSR with speech enhancement, achieving noise robustness without resorting to explicit noise mask generation. We take advantage of the bottleneck Conformer and audio reconstruction objectives to effectively refine noisy audio representations with video assistance, while preserving speech semantic integrity to facilitate subsequent cross-modal fusion. Experimental results on the LRS3 dataset show superior noise robustness compared to mask-based baseline methods, validating the efficacy of implicit noise suppression enabled by bottleneck-driven cross-modal fusion. Further work will further explore incorporating video reconstruction-based constraints for audio-visual corrupted scenarios.

### v2

@@ Abstract
Audio-visual speech recognition (AVSR) typically improves recognition accuracy in noisy environments by integrating noise-immune visual cues with audio signals. Nevertheless, high-noise audio inputs are prone to introducing adverse interference into the feature fusion process. To mitigate this, recent AVSR methods often adopt mask-based strategies to filter audio noise during feature interaction and fusion, yet such methods risk discarding semantically relevant information alongside noise. In this work, we propose an end-to-end noise-robust AVSR framework coupled with speech enhancement, eliminating the need for explicit noise mask generation. The framework leverages a Conformer-based bottleneck fusion module to implicitly refine noisy audio features with video assistance. By reducing modality redundancy and enhancing inter-modal interactions, our method aims to preserve speech semantic integrity to achieve robust recognition performance. Experimental evaluations on the public LRS3 benchmark suggest that our method outperforms prior advanced mask-based baselines under noisy conditions.
@@ 1 Introduction
Most contemporary AVSR models employ cross-modal attention mechanisms to facilitate effective audio-visual interaction [6, 7, 8, 9, 10]. Compared to early strategies such as direct feature concatenation, cross-attention achieves dynamic alignment and richer multimodal integration. But, when audio inputs are severely corrupted by noise, the interaction process is prone to disruption. Noisy audio features can introduce irrelevant or misleading information, forcing the model to undertake the entangled tasks of both implicitly denoising the input and extracting critical speech-related information simultaneously. This dual burden might overwhelm the cross-modal interaction module, leading to suboptimal feature fusion. To alleviate the negative effects of noise, recent studies have proposed custom-designed masking networks to generate noise-reduction masks, which explicitly suppress irrelevant noise components in audio features [4, 11, 12]. These methods aim to suppress noise by selectively focusing on relevant audio information during fusion. Nevertheless, they are typically driven solely by the final AVSR objective, which cannot guarantee the preservation of semantic integrity during this lossy noise-suppression process.
@@ 2 Proposed Method
Fig. 1 illustrates the proposed noise-robust AVSR framework, which takes video and noisy audio as inputs to perform speech recognition without explicit mask generation. This framework integrates a primary AVSR model with a dedicated speech enhancement (denoising) module, which generates refined audio representations from noisy features. Positioned between audio feature extraction and cross-modal fusion, the enhancement module is jointly optimized with the AVSR model. These refined audio features are subsequently fed into a cross-modal encoder, where they are fused with visual features to facilitate deep and reliable speech understanding.
@@ 2.2 Audio-Visual Bottleneck Conformer
where $\theta_{v}$ and $\theta_{a}$ are the parameters of the visual Conformer and the audio Conformer, respectively. $\|$” denotes the concatenation operation of the latent tokens. Note that the convolution module within the Conformer block is only used for modality tokens excluding the bottleneck tokens. The iterative bottleneck attention computation across $\mathbf{h}_{a}$ and $\mathbf{h}_{v}$ ultimately yields refined representations $\mathbf{z}_{v}=\mathbf{h}_{v}^{L}$ and $\mathbf{z}_{a}=\mathbf{h}_{a}^{L}$. Since all cross-modal attention flow must pass through these bottleneck tokens, such tight “fusion” bottlenecks force the model to condense modality-specific information and share only essential content. This design allows the visual modality to guide the purification of noisy audio features in a computationally efficient manner, which reduces cross-modal attention computations from $\mathcal{O}((N_{a}+N_{v})^{2})$ to $\mathcal{O}((K+N_{a})^{2})+\mathcal{O}((K+N_{v})^{2})$.
@@ 3.2 Experimental Settings
[…]
@@
[…]
@@ 3.3.1 Influence of the number of bottleneck tokens
The bottleneck tokens serve a key role in our AVBC fusion structure. We thus conduct experiments to evaluate how their quantity influences the model’s final prediction performance under -5dB babble noise. We keep the number of bottleneck tokens to be much smaller than the total latent units per modality, explicitly ensuring only necessary information resides in these bottleneck tokens. Specifically, we set the number of tokens to 0, 1, 2, 4, 8, and 16 respectively for comparison, where “0” means that the two modalities skip the bottleneck and directly engage in cross-attention. As depicted in Fig. 2, we can observe that the model delivers relatively favorable performance when the number of tokens is set to 4. The results imply that too few tokens may hinder sufficient exchange of essential inter-modal information, while too many could compromise the model’s ability to prioritize only essential content transmission.
@@ 3.3.3 Noise robustness evaluation
[…]
@@ 3.3.4 Experiments on various input conditions
Since cross-modal attention operations are decoupled from the modal encoders, our AVSR framework generalizes to single-modality input sequences, aligning with [9, 31]. To investigate the model’s robustness under diverse input conditions, we further conduct evaluations covering clean audio and overlapped speech, each assessed with and without video input. Overlapped speech is generated by randomly selecting a sample from the test set that is distinct from the input audio and overlapping it with the input audio at an SNR of -5dB. During testing, the two distinct speech signals are played simultaneously, while the input video only provides visual cues for the target speech. Notably, the bottleneck tokens within the AVBC still interact with the audio sequence even when video input is disabled. As shown in Table 3, the presence or absence of video input exerts minimal influence on the final performance under clean audio conditions. However, when overlapped speech is used as the audio input, the visual modality becomes crucial for effectively “selecting” the target speech from the mixed speech. Furthermore, our model (without speech enhancement) performs better than the baseline in [9] under overlapped speech scenarios. This advantage likely stems from the bottleneck tokens preserving modality-shared features, which serve as complementary cues for the corrupted audio.
@@ 4 Conclusion
In this paper, we propose a noise-robust AVSR architecture that integrates AVSR with speech enhancement, achieving noise robustness without resorting to explicit noise mask generation. We take advantage of the bottleneck Conformer and audio reconstruction objectives to effectively refine noisy audio representations with video assistance, while preserving speech semantic integrity to facilitate subsequent cross-modal fusion. Experimental results on the LRS3 dataset show superior noise robustness compared to mask-based baseline methods, validating the efficacy of implicit noise suppression enabled by bottleneck-driven cross-modal fusion.